A company has a policy of retiring company cars; this policy looks at number of miles driven, purpose of trips, style of car and other features. The distribution of the number of months in service for the fleet of cars is bell-shaped and has a mean of 54 months and a standard deviation of 3 months. Using the 68-95-99.7 rule, what is the approximate percentage of cars that remain in service between 60 and 63 months?
Answer #1

Solution month Girentit el=54 month, 6=3 First Standard deviation of the mean e-6=54-3=51 4+6= 54+3=57 Approximately 68% of tEmpirical Rule: 34 341 3.59 L 13.5% 2354 2.351. 2015 20151 44 47 51 54 57 60 63 2:35 % is the approximate percentage of cars